OpenStudy (anonymous):

For each of the following, determine whether the series is convergent or divergent. 1: sum x=1 to x=infinity 6/(n(n+3)) 2: sum x=1 to x=infinity ln(n)/(2n) 3: sum x=1 to x=infinity 1/(3+n) 4: sum x=1 to infinity 6/(n^2-65) 5: sum x=1 to x=infinity (1+6^n)/(3+6^n)

OpenStudy (anonymous):

It would be nice if you can also tell me which test is used to determine the convergence/divergence of the series

OpenStudy (anonymous):

You can use the integral test on the first. It suffices to show that the integral of \[\lim_{c-> \infty}\int\limits_{1}^{c}\frac{6}{x(x+3)}dx \]is convergent. When you integrate this out, you get\[\lim_{c-> \infty}[2\log (x) - 2 \log (3 + x)|_1^c=\lim_{c-> \infty}[2\log \frac{x}{3+x}|_1^c\]\[=\lim_{c-> \infty}[2\log \frac{1}{3/x+1}|_1^c=2 \log 1 - 2 \log \frac{1}{4}=4 \log 2<\infty\]

OpenStudy (anonymous):

You familiar with this test?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

It works if the terms of your series are monotonic decreasing.

OpenStudy (anonymous):

I have a feeling you could use this test for the others. The second one is divergent using the same test. Don't wast time - set it up and just let wolframalpha do the integral so you can check it out.

OpenStudy (anonymous):

The only catch is that the terms of your sequence have to be monotone decreasing.

OpenStudy (anonymous):

How about the fifth series

OpenStudy (anonymous):

I'll look at it.

OpenStudy (anonymous):

Also for the 4th one, you can't use the integral test right?

OpenStudy (anonymous):

Yeah...you don't need to use the integral test for all of them. For the fifth, you can use the limit test. The test says that if the limit as n approaches infinity of a_n does not equal zero, the series will diverge. All you have to do is show the limit of (1+6^n)/(3+6^n) does not equal zero.

OpenStudy (anonymous):

is that the nth term test?

OpenStudy (anonymous):

might be - different names

OpenStudy (anonymous):

Thanks :) How about the 4th one. Which test should i use? I just have difficult to decide what test should be used

OpenStudy (anonymous):

I'll have a look.

OpenStudy (anonymous):

sorry for bothering you :[

OpenStudy (anonymous):

You're not bothering me :)

OpenStudy (anonymous):

After a while, 6/(n^2-65) will just behave like 6/n^2 which diverges. Is that reasonable?

OpenStudy (anonymous):

I'm just going through them now - had to do something.

OpenStudy (anonymous):

:) NP

OpenStudy (anonymous):

hi guys. i have midterm from same chapter. what is your advices about which method must be use for each equation. how can i decide it ? sorry for my poor english i hope you understand it.

OpenStudy (anonymous):

I have some advice - just give me a minute.

OpenStudy (anonymous):

ok thnx (:

OpenStudy (anonymous):

I know 4 diverges, and because of its form (polynomial fraction) I'm looking into the comparison test, or the limit comparison test. I'm trying to find an appropriate divergent sequence. I have to do this because the ratio test, etc. are inconclusive, and the integral (from the integral test) does not exist over all the domain we're interested in...so hang on...a little longer...

OpenStudy (anonymous):

Thanks for your effort :)

OpenStudy (anonymous):

Okay, sorry, got distracted again...but, I think I have your answer. Look at what happens when n=5 -> the denominator does not exist. I was looking at this series and thinking, "Why is it not convergent if it's asymptotically equivalent to 1/n^2 (which is a p-series with p>1)...well, yeah, you're diving be zero at some point, which is undefined. If you exclude n=5, you'll get a convergent sum.

OpenStudy (anonymous):

Well, the denominator exists, but the quotient doesn't.

OpenStudy (anonymous):

You comfortable with that, dichalao?

OpenStudy (anonymous):

What you mean n=5 ->denominator does not exist?

OpenStudy (anonymous):

You can't have a denominator equal to 0. When n is 5, the denominator is 5^2-25 = 25 - 25 = 0. I corrected myself above: the denominator exists, it's the quotient that doesn't (at this point).

OpenStudy (anonymous):

(n^2-65) it is n^2-65

OpenStudy (anonymous):

\[\frac{1}{0}= \infty\]You can remember it loosely as that.

OpenStudy (anonymous):

and n can only be integers so ...it should be continious

OpenStudy (anonymous):

Lhunardaien, here is a good run-down/strategy sheet. http://tutorial.math.lamar.edu/Classes/CalcII/SeriesStrategy.aspx If I find any more, I'll post them here.

OpenStudy (anonymous):

This is what's happening with your series (summing the first seven terms):

OpenStudy (anonymous):

\[-\frac{1}{24}-\frac{1}{21}-\frac{1}{16}-\frac{1}{9}+\frac{1}{0}+\frac{1}{11}+\frac{1}{24}+...\]

OpenStudy (anonymous):

\[\lim_{c->0}\frac{1}{c}=\infty\]

OpenStudy (anonymous):

Assuming the rest of the series is convergent, say to limit s, then you'd have, basically,\[s+\infty=\infty\]

OpenStudy (anonymous):

It's therefore not convergent.

OpenStudy (anonymous):

For what number of n you got how u got 1/0

OpenStudy (anonymous):

you got 1/0

OpenStudy (anonymous):

\[\frac{6}{n^2-25}\rightarrow \frac{6}{5^2-25}=\frac{6}{0}=6\times \frac{1}{0}\]

OpenStudy (anonymous):

Sorry, was the 1 bit doing you in?

OpenStudy (anonymous):

Since multiplicative constants don't do much, I just leave them out whenever it's appropriate. Sorry if that added to some confusion.

OpenStudy (anonymous):

wait the original problem is (6/(n^2-65)) not 25

OpenStudy (anonymous):

Oh God...I misread it...

OpenStudy (anonymous):

It's convergent then

OpenStudy (anonymous):

LOL time to get pair of glasses :P

OpenStudy (anonymous):

At least you learnt something else!

OpenStudy (anonymous):

Let me make sure properly, though.

OpenStudy (anonymous):

Basically, it's asymptotically equivalent to 1/n^2 which is convergent (it's a p-series with p>1).

OpenStudy (anonymous):

yea haha so 1,4 converges and the rest diverges?

OpenStudy (anonymous):

so if i want to show how it is convergent. Which test should i use?

OpenStudy (anonymous):

Number 4?

OpenStudy (anonymous):

p-series test

OpenStudy (anonymous):

after acknowledging that, in the limit,\[\frac{6}{n^2-65} \]is asymptotically equivalent to \[\frac{6}{n^2}=6\frac{1}{n^2}\]

OpenStudy (anonymous):

Got It haha:) i might have one more problem would you mind help me out too

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

you need to check the other problems i didn't explicitly cover here. i didn't look through all of them.

OpenStudy (anonymous):

yea i did check the ones that you didn't cover :)

OpenStudy (anonymous):

We might think that a ball that is dropped from a height of 10 feet and rebounds to a height 7/8 of its previous height at each bounce keeps bouncing forever since it takes infinitely many bounces. This is not true! We examine this idea in this problem. A. Show that a ball dropped from a height h feet reaches the floor in {1\over4}\sqrt{h} seconds. Then use this result to find the time, in seconds, the ball has been bouncing when it hits the floor for the first, second, third and fourth times: time at first bounce = time at second bounce = time at third bounce = time at fourth bounce = B. How long, in seconds, has the ball been bouncing when it hits the floor for the nth time (find a closed form expression)? time at nth bounce = C. What is the total time that the ball bounces for? total time =

OpenStudy (anonymous):

{1\over4}\sqrt{h} = sqrt(h)/4

OpenStudy (anonymous):

I got part A..but i couldn't figure out part B

OpenStudy (anonymous):

Can you type out the expressions in the equation editor?

OpenStudy (anonymous):

We might think that a ball that is dropped from a height of 10 feet and rebounds to a height 7/8 of its previous height at each bounce keeps bouncing forever since it takes infinitely many bounces. A. Show that a ball dropped from a height h feet reaches the floor in \[{1\over4}\sqrt{h}\] seconds. Then use this result to find the time, in seconds, the ball has been bouncing when it hits the floor for the first, second, third and fourth times: time at first bounce = time at second bounce = time at third bounce = time at fourth bounce = B. How long, in seconds, has the ball been bouncing when it hits the floor for the nth time (find a closed form expression)? time at nth bounce = C. What is the total time that the ball bounces for? total time =

OpenStudy (anonymous):

Okay, I was thrown for a minute then remembered you're using imperial units.

OpenStudy (anonymous):

Haha Take your time :)

OpenStudy (anonymous):

Okay...I have an answer.

OpenStudy (anonymous):

COOL!!

OpenStudy (anonymous):

When setting up your times, you have to keep in mind, for the first time, the ball falls only to the ground (from where it's dropped) BUT AFTER THAT, the time you calculate needs to be double at each step because it spends time going back up, and then back down to mark the next turn.

OpenStudy (anonymous):

Are you expected to use mathematical induction to prove you have the correct series, or can it just be 'obvious'?

OpenStudy (anonymous):

I got part A,so i tried to come up with an expression ,,,i got I did that.. i got sqrt10/4+(sqrt10)(sqrt(7/8))^(n-1)/2 for part B...

OpenStudy (anonymous):

No just the answer is fine

OpenStudy (anonymous):

I wrote the accumulated times up to 4 for you, then skipped to s_n. I'll scan it.

OpenStudy (anonymous):

OH MY GOD you are the NICEST person i have ever met online :OOOOO

OpenStudy (anonymous):

After the line for s_n, I jumped to the limit of an infinite sum. The formula on the second page is the accumulated time for the nth step.

OpenStudy (anonymous):

Well, according to my fried head. I'm *very* tired, so please double-check everything.

OpenStudy (anonymous):

HAHA it is all good. Thanks for all your help. This really helps me out alot

OpenStudy (anonymous):

Okay. Just check the accumulated sum. I usually derive the sums of geometric series on a case-by-case basis, but this time, just pulled the formula. I don't think I'm wrong in my analysis, though, like I said, I'm fried, so the mathematics needs to be checked. Good luck.

OpenStudy (anonymous):

Lhunardaien, did you get the link?

OpenStudy (anonymous):

I am looking through it :)

OpenStudy (anonymous):

i just looked now and thanx for link im looking now.

OpenStudy (anonymous):

Lok,,can you please explain how to derive equation for the accumulated time for n step~

OpenStudy (anonymous):

Yeah, look at the powers of r in relation to the subscript on s.

OpenStudy (anonymous):

When n=2, the highest power is 1/2 n=3, the highest power is 2/2 n=4, the highest power is 3/2 ... n=k, the highest power is (k-1)/2

OpenStudy (anonymous):

brb

OpenStudy (anonymous):

\[1+r ^{1/2}+r+^{3/2}+.....) ->1(1-r^{n/2})/(1-r) this is the most confusing part

OpenStudy (anonymous):

Okay, this is just applying the formula for the sum of a geometric series. For r<1, the sum of the first n terms of the series a+ar+ar^2+...+ar^n is \[\frac{a(1-r^{n+1})}{1-r}\]

OpenStudy (anonymous):

Here, a=1, and r is the r we're using anyway.

OpenStudy (anonymous):

7/8, I think...

OpenStudy (anonymous):

I input that equation , the system indicated that the equation is incorrect

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

is this a test?

OpenStudy (anonymous):

This is a homework problem. (Webwork)

OpenStudy (anonymous):

I think your analysis makes very much sense.

OpenStudy (anonymous):

The accumulated sum is correct tho